# Sampling the space of solutions of an artificial neural network

### v1

@@ Abstract
The weight space of an artificial neural network can be systematically explored using tools from statistical mechanics. We employ a combination of a hybrid Monte Carlo algorithm which performs long exploration steps, a ratchet-based algorithm to investigate connectivity paths, and coupled replica models simulations to study subdominant flat regions. Our analysis focuses on one hidden layer networks and spans a range of energy levels and constrained density regimes.
Near the interpolation threshold, the low-energy manifold shows a spiky topology. While these spikes aid gradient descent, they can trap general sampling algorithms at low temperatures; they remain connected by complex paths in a confined, non-flat region.
In the overparameterized regime, however, the low-energy manifold becomes entirely flat, forming an extended complex structure that is easy to sample. These numerical results are supported by an analytical study of the training error landscape, and we show numerically that the qualitative features of the loss landscape are robust across different data structures. Our study aims to provide new methodological insights for developing scalable methods for large networks.
@@ I Introduction
On the empirical side, simple low-dimensional visualizations of the loss landscape have been used to gain insight into the general properties of these minima and their arrangement within the loss landscape Li et al. (2018); Huang et al. (2020). In particular, linear and piecewise linear paths Draxler et al. (2018); Garipov et al. (2018); Fort and Jastrzebski (2019) have been used to show that weight configurations found by SGD are generally not linearly connected but are nevertheless connected through a piecewise path Pittorino et al. (2022); Entezari et al. (2022).
@@
Our results are presented as follows. We shall first introduce the model and the algorithms to sample the loss space (Sect. II), then we shall describe the numerical results obtained close to the interpolation threshold (Sect. III.3) . In the case of the overparametrized regime (Sect. III.2), we shall first present the numerical results of the sampling and then compare it with the analytical calculations. Finally, we will discuss how these results change when realistic, correlated data are used as input (Sect. III.3), and we will draw the overall conclusions (Sect. IV).
@@ II.1 The model and main definitions
where $K$ is the width of the hidden layer and $\varphi$ is a non-linear activation function. The first layer is parameterized by a set of weights $\boldsymbol{w}\in\mathbb{R}^{N}$, which will be trained to learn a dataset. The weights of the second layer, $c_{l}$ are fixed (not learned) to ±$\pm 1$, with equal probability. Thus, the number of learned weights is exactly equal to $N$, the input dimension of the network.
@@
We consider a syntetic dataset $\mathcal{D}$ made of $P$ input vectors $\boldsymbol{x}^{\mu}$, whose elements are extracted from a standard normal distribution, and the associated labels $y^{\mu}=\pm 1$, also generated randomly with equal probability. The ratio $\alpha=P/N$ is called constrained density.
@@
is equal to zero, where we have denoted by $\Theta(\cdot)$ the Heaviside function, and by $\ell_{\mathrm{NE}}$ the “error counting” loss. Since this loss is not differentiable, it is not used for optimization in machine learning, which typically relies on gradient-based algorithms. Instead, the error counting loss usually serves to evaluate whether a solution has been found, rather than guiding the optimization process itself.
@@
Indeed in the limit $\beta\to\infty$ the Boltzmann measure of Eq. (7), equipped with the loss function of Eq. (6), focuses on solutions satisfying Eq. (5). We stress that extracting solutions by optimizing a loss different from the error counting loss is therefore to be considered as "atypical".
@@ II.2 The canonical–ensemble framework
where $\beta=1/T$ is the inverse temperature, $p(\boldsymbol{w})$ is the prior distribution of the weights. For simplicity, a Boltzmann constant $k_{B}=1$ is assumed in every equation. The factor $Z(\beta;\mathcal{D})$ is a normalization factor that is called evidence in Bayesian statistics and partition function in statistical physics and it reads
@@
The loss function usually considered in the machine learning literature is factorized over the $P$ elements of the dataset, i.e.
@@
where $\ell$ is a loss function per pattern and $\Delta(\boldsymbol{w})$ identifies the preactivation of the output node. We consider as a prior a standard normal distribution, or equivalently a $L^{2}$ regularization term with parameter $\lambda$
@@ II.3 Hybrid Monte Carlo algorithm
Operatively, at each iteration of the Metropolis algorithm, the momenta are extracted from a Maxwell–Boltzmann distribution at temperature $T$ and a trajectory in the phase space is generated solving Eqs. (15). The final point of the trajectory is then accepted with probability $\min(1,\exp[-\Delta H/T])$, where $\Delta H=H(\boldsymbol{w}_{i+1},\boldsymbol{p}_{i+1})-H(\boldsymbol{w}_{i},\boldsymbol{p}_{i})$. In the condition of detailed balance, the kinetic energy arising from the Metropolis acceptance probability and that coming from the Maxwell–Boltzmann term in the a priori probability $p_{ap}$ cancel out and the algorithm converges to Eq. (12).
The power of this scheme is that it can produce large moves in a direction that change smoothly the energy of the system, and thus the Metropolis acceptance rate can remain high.
@@ II.4 Double ratchet
In deep learning, characterizing the structure of neural network loss landscapes is crucial for understanding both optimization dynamics and generalization properties. A key line of research explores pathways between distinct weight configurations $\boldsymbol{w}$ and $\boldsymbol{w}^{\prime}$, which typically correspond to solutions with low loss. The simplest approach examines how the loss behaves along the linear (or straight) path connecting these points. As shown in Draxler et al. (2018), overparameterized neural networks trained with stochastic gradient descent (SGD) often exhibit a loss barrier along the linear path. The barrier can be notably decreased by removing the symmetries that the neural network possess, like permutation symmetry of the hidden units Pittorino et al. (2022); Entezari et al. (2022).
@@
$U_{rt}\bigl{[}\boldsymbol{w}_{1}(t),\boldsymbol{w}_{2}(t)\bigr{]}\\ =U\bigl{[}\boldsymbol{w}_{1}(t)\bigr{]}+U\bigl{[}\boldsymbol{w}_{2}(t)\bigr{]}+\widetilde{U}\bigl{[}\boldsymbol{w}_{1}(t),\boldsymbol{w}_{2}(t)\bigr{]}$ (17)
@@
$\widetilde{U}\bigl{[}\boldsymbol{w}_{1}(t),\boldsymbol{w}_{2}(t)\bigr{]}=\begin{cases}\frac{k}{2}\bigl{[}|\boldsymbol{w}_{1}(t)-\boldsymbol{w}_{2}(t)|-d_{\min}(t)\bigr{]}^{2}&\text{if }|\boldsymbol{w}_{1}(t)-\boldsymbol{w}_{2}(t)|>d_{\mathrm{min}}(t)\\ 0&\text{if }|\boldsymbol{w}_{1}(t)-\boldsymbol{w}_{2}(t)|\leq d_{\min}(t)\end{cases}$ (18)
@@
During each simulation, the two vectors are updated sequentially at each time step according to Eq. (17), until their cosine–similarity (or normalized overlap)
@@
is equal to 1. We will call the point $\boldsymbol{s}^{\star}$ where the two trajectories $\boldsymbol{w}_{1}(t)$ and $\boldsymbol{w}_{2}(t)$ meet as the “anchor” weight of $\boldsymbol{w}_{1}$, $\boldsymbol{w}_{2}$.
@@ II.5 Coupled replica simulations
Algorithmically, the weights $\{\boldsymbol{w}_{i}\}_{i=1}^{y}$ are updated sequentially according to Eq. (21), whereas the barycenter is computed once one of the replicas has moved. The value of $\gamma$ is increased during the simulations, until the $y$ replicas $\boldsymbol{w}_{i}$ all collapse onto a single high–entropy weight $\boldsymbol{c}$, that we will refer in the rest of the paper as the “center” weight found by the coupled replica simulation.
@@ III Results
Unless stated otherwise, in all the numerical simulations we have used a tree committee machine with $N=1000$ input neurons, $K=50$ hidden neurons, and ReLU activation functions $\varphi(x)=\max(0,x)$. In HMC simulations we have used the cross entropy loss function in Eq. (4). The norm of the weights is controlled by the Lagrange multiplier $\lambda=2P\cdot 10^{-7}$, with $P$ being the dimension of the dataset. The mass matrix in Eq. (14) required by the algorithm is set $\mathcal{M}=\mathbb{I}$.
Figure 2: a) The average energy $\langle U\rangle_{\boldsymbol{s}}$ (blue dots) and the mean training error $\langle\epsilon\rangle_{\boldsymbol{s}}$ (orange dots) computed on GD solutions with respect to the constraint density $\alpha=P/N$. b) The average energy $\langle U\rangle_{T}$ (blue dots) and the mean training error $\langle\epsilon\rangle_{T}$ (orange dots) with respect to the temperature $T$. The gray vertical lines identify the three temperatures regimes (low, medium and high). c) The cosine–similarity distribution among solutions $\boldsymbol{s}$ at $\alpha=1.8$ (upper panel, in green). Middle panel, in blue: the intra (thin line histograms) and the inter (thick line histogram) state overlap distribution $\rho(q)$ at low–temperature ($T=1.8\cdot 10^{-2}$). Lower panel, in orange: the intra and inter state overlap distribution $\rho(q)$ at intermediate temperature ($T=1.8\cdot 10^{1}$, lower panel, in orange). The three overlapping distributions have been slightly shifted along the $x$–axis to facilitate the comparison. d) A sketch of the subspace of solutions at fixed norm as suggested by the HMC simulations.
@@ III.1 The model close to the interpolation threshold
We first study the loss landscape of the model in the underparametrized regime, i.e. just below the threshold at which full–batch GD can no longer find weights with zero training error and which is usually called "interpolation threshold" Engel and Van den Broeck (2001); Belkin et al. (2019). This happens approximately around $\alpha=1.8$ (Fig. 2a), which is very far from the SAT/UNSAT transition $\alpha_{c}\simeq 2.65$ where solutions cease to exist and which was computed in Annesi et al. (2024). All the numerical studies of this section therefore refer to the case $\alpha=1.8$.
@@ III.1.1 The weight space displays three regimes with respect to the temperature
In the high temperature phase, the average energy deviates from the linear behavior and the training error further increases. The distribution of similarities $q$ is very broaden. However, here the HMC algorithm becomes inefficient because the high velocities extracted before each HMC move require a very small step $\delta t$ for a correct integration of the equations of motion and thus for an acceptable acceptance rate.
@@ III.1.2 The low–energy basins are connected by complex paths
The double ratchets are initialized on different pairs of solutions $(\boldsymbol{s}_{i},\boldsymbol{s}_{j})$ found after a short thermalization of the system of approximately $10^{4}$ HMC steps. Along each simulation, the weights do not cross significant energy barriers, as compared to the average energy at this temperature, keeping the training error equal to zero. Conversely, the linear interpolation between pairs of points along the double–ratchet trajectory results in barriers that are substantially higher than the average energy (Fig. 3b). These results suggest that, although the solutions found by gradient descent are linearly mode disconnected, low-energy tortuous paths joining them still exist.
The anchor point $\boldsymbol{s}^{\star}$ obtained at the end of the double–ratchet trajectory has an energy that is slightly lower than the one of the initial points $(\boldsymbol{s}_{i},\boldsymbol{s}_{j})$ and $\langle U\rangle_{\boldsymbol{s}^{\star}}=(6.80\pm 0.09)\cdot 10^{-2}$ is compatible with the average $\langle U\rangle_{T_{L}}$, that is the constrained dynamic arising from the double–ratchet simulation helps the equilibration of the system. The anchor weights $\boldsymbol{s}^{\star}$ obtained from several double ratchet simulations initialized from different weight pairs $(\boldsymbol{s}_{i},\boldsymbol{s}_{j})$ exhibit larger similarity than the respective initial vectors and are separated by energy barriers along their linear interpolation (Fig. 3c). This suggests the presence of a compact and narrow region of the solution space that must be traversed to move from one gradient descent solution to another. The similarity between the anchor weights $\boldsymbol{s}^{\star}$ and the associated initial weights is also centered around $0.6$, as that between the starting solutions.

### v2

@@ Abstract
The weight space of an artificial neural network can be systematically explored using tools from statistical mechanics. We employ a combination of a hybrid Monte Carlo algorithm which performs long exploration steps, a ratchet-based algorithm to investigate connectivity paths, and coupled replica models simulations to study subdominant flat regions. Our analysis focuses on one hidden layer networks and spans a range of energy levels and constrained density regimes. Near the interpolation threshold, the low-energy manifold shows a spiky topology. In the overparameterized regime, however, the low-energy manifold becomes entirely flat, forming an extended complex structure that is easy to sample. These numerical results are supported by an analytical study of the training error landscape, and we show numerically that the qualitative features of the loss landscape are robust across different data structures. Our study aims to provide new methodological insights for developing scalable methods for large networks.
@@ I Introduction
On the empirical side, simple low-dimensional visualizations of the loss landscape have been used to gain insight into the general properties of these minima and their arrangement within the loss landscape Li et al. (2018a); Huang et al. (2020). In particular, linear and piecewise linear paths Draxler et al. (2018); Garipov et al. (2018); Fort and Jastrzebski (2019) have been used to show that weight configurations found by SGD are generally not linearly connected but are nevertheless connected through a piecewise path Pittorino et al. (2022); Entezari et al. (2022).
@@
A substantial body of work exists on the theoretical analysis of single-hidden-layer neural networks with fully connected receptive fields in the first layer, both when trained to store random patterns Engel et al. (1992); Urbanczik (1997); Nishiyama and Ohzeki (2025) or when the label is given by a teacher with a similar architecture Schwarze and Hertz (1992); Schwarze (1993); Schwarze and Hertz (1993a, b); Aubin et al. (2018); Oostwal et al. (2021); Citton et al. (2025); Barbier et al. (2025a). Recently, the scaling regime $K=O(N)$ has also been examined in Barbier et al. (2025b).
Our results are presented as follows. We shall first introduce the model and the algorithms to sample the loss space (Sect. II), then we shall describe the numerical results obtained close to the interpolation threshold (Sect. III.1) . In the case of the overparametrized regime (Sect. III.2), we shall first present the numerical results of the sampling and then compare it with the analytical calculations. Finally, we will discuss how these results change when realistic, correlated data are used as input (Sect. III.3), and we will draw the overall conclusions (Sect. IV).
@@ II.1 The model and main definitions
where $K$ is the width of the hidden layer and $\varphi$ is a non-linear activation function. The first layer is parameterized by a set of weights $\boldsymbol{w}\in\mathbb{R}^{N}$, which will be trained to learn a dataset. The weights of the second layer, $c_{l}$ are fixed to ±$\pm 1$, with equal probability. Thus, the number of learned weights is exactly equal to $N$, the input dimension of the network. The main reason for not learning the weights of the second layer is to eliminate the symmetries of node permutations, which would make analyzing the space of solutions more difficult.
@@
We consider a synthetic dataset $\mathcal{D}$ made of $P$ input vectors $\boldsymbol{x}^{\mu}$, whose elements are extracted from a standard normal distribution, and the associated labels $y^{\mu}=\pm 1$, also generated randomly with equal probability. The ratio $\alpha=P/N$ is called constrained density.
@@
is equal to zero, where we have denoted by $\Theta(\cdot)$ the Heaviside function, and by $\ell_{\mathrm{NE}}$ the “error counting” (or number of error) loss. Since this loss is not differentiable, it is not used for optimization in machine learning, which typically relies on gradient-based algorithms. The loss that is used to find solutions is usually called “training loss” in the machine learning jargon. Instead, the error counting loss usually serves to evaluate whether a solution has been found [^1] In general the training error $\epsilon_{t}$ could be zero even if the training loss is not., rather than guiding the optimization process itself.
@@
We stress that extracting solutions by optimizing a loss different from the error counting loss is therefore to be considered as “atypical”.
@@ II.2 The canonical–ensemble framework
where $\beta=1/T$ is the inverse temperature, $p(\boldsymbol{w})$ is the prior distribution of the weights. For simplicity, a Boltzmann constant $k_{B}=1$ is assumed in every equation. The factor $Z(\beta;\mathcal{D})$ is the partition function and it is equal (up to a normalization factor) to what is called evidence in Bayesian statistics; it reads
@@
The loss function usually considered in the machine learning literature is separable into a sum over the $P$ elements of the dataset, i.e.
@@
where $\ell$ is a loss function per pattern and $\Delta(\boldsymbol{w})$ identifies the preactivation of the output node. A solution satisfying Eq. (5) can be obtained from the canonical-ensemble framework by performing the limit $\beta\to\infty$ of the Boltzmann measure in Eq. (7), equipped with the loss function of Eq. (6).
In the following we consider as a prior a standard normal distribution, or equivalently a $L^{2}$ regularization term with parameter $\lambda$
@@ II.3 Hybrid Monte Carlo algorithm
Operatively, at each iteration of the Metropolis algorithm, the momenta are extracted from a Maxwell–Boltzmann distribution at temperature $T$ and a trajectory in the phase space is generated solving Eqs. (15). The final point of the trajectory is then accepted with probability $\min(1,\exp[-\Delta H/T])$, where $\Delta H=H(\boldsymbol{w}_{i+1},\boldsymbol{p}_{i+1})-H(\boldsymbol{w}_{i},\boldsymbol{p}_{i})$. In the condition of detailed balance (i.e. when a time–reversible integrator is employed), the kinetic energy arising from the Metropolis acceptance probability and that coming from the Maxwell–Boltzmann term associated with the extraction of the random momenta cancel out. This leads the distribution of the states sampled by HMC to converge to Eq. (12).
The power of this scheme is that it can produce large moves in a direction that smoothly changes the energy of the system, and thus the Metropolis acceptance rate can remain high.
@@ II.4 Double ratchet
A key line of research in characterizing the structure of neural network loss landscapes explores pathways between distinct weight configurations $\boldsymbol{w}$ and $\boldsymbol{w}^{\prime}$, which typically correspond to solutions with low loss Goodfellow et al. (2014); Garipov et al. (2018); Fort and Jastrzebski (2019); Li et al. (2018b); Frankle (2020). The simplest approach examines how the loss behaves along the linear (or straight) path connecting these points. As shown in Draxler et al. (2018), overparameterized neural networks trained with stochastic gradient descent (SGD) often exhibit a loss barrier along the linear path. The barrier can be notably decreased by removing the symmetries that the neural network possess, like permutation symmetry of the hidden units Pittorino et al. (2022); Entezari et al. (2022).
@@
$U_{rt}\bigl{[}\boldsymbol{w}_{1},\boldsymbol{w}_{2},t\bigr{]}=U\bigl{[}\boldsymbol{w}_{1}\bigr{]}+U\bigl{[}\boldsymbol{w}_{2}\bigr{]}+\widetilde{U}\bigl{[}\boldsymbol{w}_{1},\boldsymbol{w}_{2},t\bigr{]}$ (17)
@@
$\widetilde{U}\bigl{[}\boldsymbol{w}_{1}(t),\boldsymbol{w}_{2}(t),t\bigr{]}=\begin{cases}\frac{k}{2}\bigl{[}|\boldsymbol{w}_{1}(t)-\boldsymbol{w}_{2}(t)|-d_{\min}(t)\bigr{]}^{2}&\text{if }|\boldsymbol{w}_{1}(t)-\boldsymbol{w}_{2}(t)|>d_{\mathrm{min}}(t)\\ 0&\text{if }|\boldsymbol{w}_{1}(t)-\boldsymbol{w}_{2}(t)|\leq d_{\min}(t)\end{cases}$ (18)
@@
A suitable range of values for the harmonic constant k can easily be estimated by trial and error. At high k, one of the two vectors becomes trapped along suboptimal paths, resulting in a lower acceptance rate as the system becomes stuck between the steep parabolic potential and the energy barrier separating it from the other vector. In these cases, reducing the steepness of the parabolic potential allows for larger oscillations in the distance between the two weights.
In practice, each iteration consists of two sequential updates, one for each vector. During each update, the HMC scheme uses the potential $U_{rt}\bigl{[}\boldsymbol{w}_{1}(t),\boldsymbol{w}_{2}(t)\bigr{]}$ is used for the selected vector, while the other vector is kept fixed. This process is repeated until the cosine similarity (or normalized overlap) between the two weights
@@
is $>0.995$ (i.e., $q\approx 1$). We will call the point $\boldsymbol{s}^{\star}$ where the two trajectories $\boldsymbol{w}_{1}(t)$ and $\boldsymbol{w}_{2}(t)$ meet as the “anchor” weight of $\boldsymbol{w}_{1}$, $\boldsymbol{w}_{2}$.
@@ II.5 Coupled replica simulations
The implementation of the replica simulations is similar to that of the double-ratchet. Each iteration consists of a number of sequential updates, equal to the number of replicas. During each update, the HMC scheme with potential $U_{rp}(\{\boldsymbol{w}_{i}(t)\}_{i=1}^{y})$ is adopted for the selected replica, while the other ones are kept fixed. The barycenter vector is also updated, both during the proposed move and at the end of the acceptance procedure, for each replica. The value of $\gamma$ is increased during the simulations, until the $y$ replicas $\boldsymbol{w}_{i}$ all collapse onto a single high–entropy weight $\boldsymbol{c}$, that we will refer in the rest of the paper as the “center” weight found by the coupled replica simulation.
@@ III Results
Unless stated otherwise, in all the numerical simulations we have used a tree committee machine with $N=1000$ input neurons, $K=50$ hidden neurons, and ReLU activation functions $\varphi(x)=\max(0,x)$. In HMC simulations we have used the cross entropy loss function in Eq. (4). The norm of the weights is controlled by the Lagrange multiplier $\lambda=2P\cdot 10^{-7}$, with $P=\alpha N$ being the dimension of the dataset. The mass matrix in Eq. (14) required by the algorithm is set $\mathcal{M}=\mathbb{I}$.
Figure 2: a) The average energy $\langle U\rangle_{\boldsymbol{s}}$ (blue dots) and the mean training error $\langle\epsilon\rangle_{\boldsymbol{s}}$ (orange dots) computed on GD solutions with respect to the constraint density $\alpha=P/N$. b) The average energy $\langle U\rangle_{T}$ (blue dots) and the mean training error $\langle\epsilon\rangle_{T}$ (orange dots) with respect to the temperature $T$. The gray vertical lines identify the three temperatures regimes (low, medium and high). c) The upper panel shows the cosine–similarity distribution among GD solutions $\boldsymbol{s}$ at $\alpha=1.8$ (in green). The thin–line histogram in the middle panel (in blue) shows the two intra-state overlap distributions $\rho(q)$ in the low–temperature ($T=1.8\cdot 10^{-2}$) simulations starting from two different initial conditions. The thick line histogram stands for the corresponding inter-state distribution computed between the two sets of samples states. Two thin–line histograms in the lower panel (in orange) shows intra-state overlap distributions $\rho(q)$ relative to two low–temperature ($T=1.8\cdot 10^{1}$) simulations starting from different initial conditions. The thick line histogram stands for the corresponding inter-state distribution computed between the two sets of samples states. In the lower panel, the three overlapping distributions have been slightly shifted along the $x$–axis to facilitate the comparison. d) A sketch of the subspace of solutions at fixed norm as suggested by the HMC simulations.
@@ III.1 The model close to the interpolation threshold
We first study the loss landscape of the model in the underparametrized regime, i.e. just below the threshold at which full–batch GD can no longer find weights with zero training error and which is usually called "interpolation threshold" Engel and Van den Broeck (2001); Belkin et al. (2019). This happens approximately around $\alpha=1.8$ (Fig. 2a), which is very far from the SAT/UNSAT transition $\alpha_{c}\simeq 2.65$ where solutions cease to exist and which was computed in Annesi et al. (2025). All the numerical studies of this section therefore refer to the case $\alpha=1.8$ (i.e. $P=1800$). The results below are obtained from different simulations, with the same dataset, which have reached a stationary regime. Given the dimension $N$ of the network and the dataset size $P$, we believe that these results are not instance-dependent.
@@ III.1.1 The weight space displays three regimes with respect to the temperature
In the high temperature phase, the average energy deviates from the linear behavior and the training error further increases. The distribution of similarities $q$ is very broad. However, here the HMC algorithm becomes inefficient because the high velocities extracted before each HMC move require a very small step $\delta t$ for a correct integration of the equations of motion and thus for an acceptable acceptance rate.
@@ III.1.2 The low–energy basins are connected by complex paths
The double ratchets are initialized on different pairs of solutions $(\boldsymbol{s}_{i},\boldsymbol{s}_{j})$ found after a short thermalization of the system of approximately $10^{4}$ HMC steps. Along each simulation, the weights do not cross significant energy barriers, as compared to the average energy at this temperature, keeping the training error equal to zero. Conversely, the geodesic between pairs of points (i.e. the linear interpolation between weights at the fixed value of the norm imposed by the two endpoints) along the double–ratchet trajectory results in barriers that are substantially higher than the average energy (Fig. 3b). These results suggest that, although the solutions found by gradient descent are linearly mode disconnected, low-energy tortuous paths joining them still exist.
The anchor point $\boldsymbol{s}^{\star}$ obtained at the end of the double–ratchet trajectory has an energy that is slightly lower than the one of the initial points $(\boldsymbol{s}_{i},\boldsymbol{s}_{j})$ and $\langle U\rangle_{\boldsymbol{s}^{\star}}=(6.80\pm 0.09)\cdot 10^{-2}$ is compatible with the average potential energy at $T_{L}=1.8\cdot 10^{-2}$, that is the constrained dynamic arising from the double–ratchet simulation helps the equilibration of the system. The anchor weights $\boldsymbol{s}^{\star}$ obtained from several double ratchet simulations initialized from different weight pairs $(\boldsymbol{s}_{i},\boldsymbol{s}_{j})$ exhibit larger similarity than the respective initial vectors and are separated by energy barriers along their linear interpolation (Fig. 3c). This suggests the presence of a compact and narrow region of the solution space that must be traversed to move from one gradient descent solution to another. The similarity between the anchor weights $\boldsymbol{s}^{\star}$ and the associated initial weights is also centered around $0.6$, as that between the starting solutions.